solve the floowing radical
sq rt x+3 = x-3

Answer by checkley77(12844)

sqrtx+3=x-3
X+3=(X-3)^2
X+3=X^2-6X+9
X^2-6X-X+9-3=-0
X^2-7+6=0
(X-6)(X-1)=0
X-6=0
X=6 ANSWER.
X-1=0
X=1 ANSWER.